# Tianjun Gu1 & Linfeng Li1 & Xuhong Wang2 & Chenghua Gong1 & Jingyu Gong1 & Zhizhong Zhang1 & Yuan Xie1,3 & Lizhuang Ma1 & Xin Tan1,2

### v3

@@ 3.2 Ventral Stream
The Ventral Stream, analogous to the “what” pathway in human cognition, integrates two key components: RAG-VLM (Section 3.2.1) for semantic knowledge processing and Policy-VLM (Section 3.2.2) for decision-making.
3.2.1 RAG-VLM
To build a comprehensive understanding of the task, RAG-VLM leverages the extensive world knowledge embedded within a vision-language model. Upon receiving the task $T$, the system extracts key semantic attributes (general description, appearance features, structure/shape, and common location) to efficiently build the task database.
Figure 4: The structure RAG-VLM in Ventral Stream, handling the task “NAVIGATE TO THE NEAREST SOFA”.
Figure 4 illustrates this knowledge extraction process. The structured information gathered across these dimensions forms a database of the task $T$. This representation enables the agent not only to verify whether an object encountered during navigation matches the task description but also to potentially inform planning by suggesting likely areas to explore first, thereby interfacing with the spatial reasoning components of the Dorsal Stream.
3.2.2 Policy-VLM
The Policy-VLM combines visual observations, spatial awareness, and task semantics to determine optimal actions. It utilizes the reasoning capabilities of large vision-language models through Chain-of-Thought (CoT). The CoT breaks down the complex navigation task into interpretable sub-steps: current state analysis, memory integration, goal analysis, scene assessment, path planning, and action decision.
@@ 3.3 Nav-Ensurance
To enhance the evaluation of safety and efficiency in navigation, we present a new metric, the Area Overlap Redundancy Index (AORI) (Section 3.3.1). Additionally, we develop Nav-Ensurance, including Multimodal Stuck Detection (Section 3.3.2), context-aware escape strategies (Section 3.3.3), and adaptive precision navigation (Section 3.3.4), to ensure that navigation systems operate reliably and effectively.
@@ 3.3.1 Area Overlap Redundancy Index (AORI)
where $r_{\text{overlap}}$ represents the ratio of revisited areas to total observed areas, $d_{\text{norm}}$ is the normalized density, and $w_{c}=0.8,w_{d}=0.2$ are weighting coefficients. For further details, refer to the Appendix D.
@@ 3.3.2 Multimodal Stuck Detection
Agent detects navigation stuck by analyzing metrics calculated over a sliding window of step $T$:
@@
A weighted scoring function combines these metrics:
$S=w_{\eta}\cdot\mathbb{I}[\eta<\tau_{\eta}]+w_{\rho}\cdot\mathbb{I}[\rho>\tau_{\rho}].$ (6)
Agent confirms stuck when $S\geq S_{\text{th}}$ persists for $k$ consecutive windows. This formula detects whether the agent is stuck or spinning during the navigation process through $\eta$ and $\rho$.
@@ 3.3.3 Context-aware Escape Strategies
When a stuck state is detected, the system selects an appropriate escape strategy based on the perceived information from Dorsal Stream(Section 3.1). For instance, in corner traps (perceived dead ends), a large turn (near 180) is executed. In narrow passages, a small backward step followed by a randomized direction change is employed. If the environmental context is ambiguous, the agent will analyze recent successful movement directions and attempt to move perpendicularly, significantly improving escape capabilities from complex trap situations.
@@ 3.3.4 Adaptive Precision Navigation
Additionally, when activating the precision navigation mode, the system can utilize visual analysis (using VLM) to create more detailed action options, thereby maximizing final positioning accuracy relative to the task.
@@ 4 Experiments
We evaluate our proposed DORAEMON within the Habitat simulator [32] on three large-scale datasets: HM3Dv1 (using HM3D-Semantics-v0.1 [29] from the 2022 Habitat Challenge, featuring 2000 episodes across 20 scenes with 6 goal categories), HM3Dv2 (using HM3D-Semantics-v0.2 [40] from the 2023 Habitat Challenge, with 1000 episodes across 36 scenes and 6 goal categories), and MP3D [8] from the 2021 Habitat Challenge, comprising 2195 episodes across 11 scenes with 21 goal categories. We also include evaluations on GOAT [18] (using HM3D-Semantics-v0.2), a benchmark focusing on generalized object semantics with 1000 validation episodes across 100 scenes and 25 object categories.
@@
The action space includes a stop action, a move_forward action where the distance parameter is sampled from the continuous range $[0.5\text{m},1.7\text{m}]$, and a rotate action. We adopt standard metrics to evaluate navigation performance: Success Rate (SR), the percentage of episodes where the agent successfully stops near a target object; Success weighted by Path Length (SPL), defined as $\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}}{\max(p_{i},l_{i})}$, rewarding both success and efficiency; and our proposed Area Overlap Redundancy Index (AORI) (Equation (4)), which quantifies navigation intelligence by penalizing redundant exploration (lower is better). More information is set in the Appendix E.
@@
We compare DORAEMON against several state-of-the-art object navigation methods on the HM3Dv2[40], HM3Dv1[29], and MP3D[8]. Our main comparison focuses on End-to-End Vision-Language Model (VLM) approaches [25, 14]. Beyond these direct End-to-End counterparts, we also consider a broader set of recent methods for non-End-to-End object navigation methods. More baseline details are set in the Appendix I.
@@ 4.1 Methods Comparision
Comprehensive Methods Analysis:
To ensure a fair comparison with the above non-End-to-End methods that utilize a discrete action set $\mathcal{A}$: move forward 0.25m, turn left/turn right 30, look up/lookdown 30, stop, and a common 500 steps episode limit, we conduct an additional set of experiments. In these, we normalize our agent’s interactions by approximating an equivalent number of standard discrete steps for each of DORAEMON’s actions.
Compared to the non-End-to-End approach in the Table 2, DORAEMON achieves state-of-the-art performance on SR, despite normalizing our action to set $\mathcal{A}$. Each action performed by ours corresponds to several actions in this set (details are provided in the Appendix B).
@@
DORAEMON (Ours) $\checkmark$ $\checkmark$ $\checkmark$ 55.6 21.4 66.5 20.6 41.1 15.8
Ablation Studies:
1) The effect of different modules: To represent the contribution of each module, we compared three variants (Dorsal Stream, RAG-VLM of Ventral Stream, and Policy-VLM of Ventral Stream) on HM3D v2. Removing the Dorsal Stream and RAG-VLM implies that the model relies solely on the Policy-VLM of the Dorsal Stream in decision-making. The results reported for SR, SPL, and AORI, as presented in Table 3(a), respectively, show the effectiveness of Dorsal Stream and Ventral Stream. 2) We further evaluated the performance of different Visual Language Models (VLMs), as shown in Table 3(b). The Gemini-1.5-Pro demonstrated outstanding capabilities in this task. Even when using a smaller model, our approach yielded excellent results, indicating that our framework is effective on its own rather than solely relying on the reasoning capabilities of the VLMs. More importantly, as VLMs continue to evolve, the effectiveness of our plug-and-play approach will also improve.
Table 3: Ablation of HM3Dv2 (100 episodes)
(a) Ablation of different modules
Method SR (%) $\uparrow$ SPL (%) $\uparrow$ AORI (%) $\downarrow$
w/o Dorsal & Ventral Stream 51.6 18.3 61.5
w/o Dorsal & RAG-VLM 54.0 19.8 59.1
w/o Dorsal Stream 59.0 22.7 56.3
w/o Nav-Ensurance 60.0 22.5 54.9
DORAEMON 61.0 23.7 48.8
(b) Ablation of different VLMs
VLM SR (%) $\uparrow$ SPL (%) $\uparrow$ AORI (%) $\downarrow$
Qwen-7B 49.5 20.6 68.7
Gemini-1.5-Flash 58.0 20.1 54.8
Gemini-2-Flash 59.0 21.5 57.9
Gemini-1.5-Pro 61.0 23.7 48.8
@@ 5 Conclusion
In this paper, we present DORAEMON (Decentralized Ontology-aware Reliable Agent with Enhanced Memory Oriented Navigation), a novel cognitive-inspired framework consisting of Ventral and Dorsal Streams that mimics human navigation capabilities. The Dorsal Stream implements the Hierarchical Semantic-Spatial Fusion and Topology Map to handle spatiotemporal discontinuities, while the Ventral Stream combines RAG-VLM and Policy-VLM to improve decision-making. Our approach also develops Nav-Ensurance to ensure navigation safety and efficiency. Extensive experimental results demonstrate the superior performance of DORAEMON.
@@ Appendix A Action Proposer
DORAEMON employs an Action Proposer[15] to generate a refined set of candidate actions, which the Policy-VLM then evaluates for the final action decision. As shown in Figure 5, first parameterized action candidates $A_{\text{init}}^{t}$ are generated by the parameterized action space (Equation (8)). Second, adaptive filtering (Equation (9)) refines $A_{\text{cand}}^{t}$ using exploration state $\mathcal{V}_{t}$ and historical patterns $\mathcal{H}_{t}$. Safety-critical recovery (Equation(10)) enforces a rotation cooldown $\gamma$ through viability evaluation $\mathcal{F}(\cdot)$. Finally, the projection module visually encodes $A_{\text{final}}^{t}$ into $I^{t}_{\text{anno}}$ with numeric tagging (0 for rotation) to interface with VLM’s semantic space.
Figure 5: Action proposal: (a) Collision-free action generation within ±$\pm\theta_{\text{max}}$ FOV, (b) Exploration-aware filtering with $\Delta\theta$ angular resolution, (c) Safety-constrained and action projection.
@@ Appendix B Steps Conversion
$T_{\text{episode}}=\sum_{t=1}^{500}N_{t}\leq 500$ (11)
where $N_{t}$ denotes converted steps for action at time step $t$. During our experiments, one DORAEMON step $t$ was equivalent to about 7-8 $N_{t}$
@@
4:    return $1$triangleright Explicit stop handling
@@
6:    $s_{r}\leftarrow\lceil r/\Delta_{r}\rceil$triangleright Radial step calculation
7:    $\theta_{\text{deg}}\leftarrow 180|\theta|/\pi$triangleright Radian-degree conversion
8:    $s_{\theta}\leftarrow\lceil\theta_{\text{deg}}/\Delta_{\theta}\rceil$triangleright Angular step calculation
9:    $N\leftarrow\max(s_{r}+s_{\theta},1)$triangleright Step composition
10:    return $N$
@@ Appendix C Navigation Case
Figure 7 depicts a full navigation episode in which our memory-augmented agent searches for a chair in an unfamiliar indoor scene. The seven consecutive frames show the agent’s visual observations and planned motions from entry to target acquisition. Throughout the sequence, the agent (i) reasons about semantic priors—chairs are more likely near tables or in living/dining areas; (ii) fuses transient visual evidence with its episodic memory to avoid revisiting explored regions; and (iii) selects actions that maximise expected information gain while respecting safety constraints. The case, therefore, illustrates how the proposed memory module complements on-the-fly perception to yield efficient, goal-directed exploration in cluttered, real-world layouts.
@@
Figure 7: Navigation case Each row shows one decision step. Left: the green circle highlights the action selected for this step. Upper-right dashed green box displays the most relevant episodic memory retrieved at this step. Lower-right speech bubble is the agent’s natural-language rationale that fuses (i) semantic priors, (ii) current visual evidence, and (iii) memory cues.
@@ Appendix D Detailed Description of AORI
D.0.1 Area Overlap Redundancy Index (AORI)
@@
  • Map resolution: $000\times 5,$ 000 × 5 , 000 grid (map_size=5000)
@@
  • Optimal Case (AORI=0): When $\text{When }r_{\text{overlap}}=0\ \&\ d_{\text{normalized}}=0\Rightarrow 1-(0.8\cdot 1^{2}+0.2\cdot 1)=0$
  • Worst Case (AORI=1):When $\text{When }r_{\text{overlap}}=1\ \&\ d_{\text{normalized}}=1\Rightarrow 1-(0.8\cdot 0+0.2\cdot 0)=1$
@@
    $r_{\text{overlap}}$ $=\frac{99}{99}=1.0,$ (15)
    $\lambda$ $=0.8\cdot\frac{\pi(60/5000)^{2}}{1}\cdot 100\approx 0.014,$
    $d_{\text{norm}}$ $=\min\left(1,\frac{100}{0.014}\right)=1.0,$
    AORI $=1-[0.8(1-1)^{2}+0.2(1-1)]=1.0$
@@
    $r_{\text{overlap}}$ $\approx\frac{38}{499}\approx 0.076,$ (16)
    $\lambda$ $=0.8\cdot\frac{\pi(60/5000)^{2}}{1}\cdot 500\approx 0.069,$
    $d_{\text{norm}}$ $=\min\left(1,\frac{62}{0.069}\right)=1.0,$
    AORI $=1-[0.8*(1-0.076)^{2}+0.2*(1-1)]\approx 0.285$

### v4

@@ 3.2 Ventral Stream
The Ventral Stream, analogous to the “what” pathway in human cognition, integrates two key components: CoDe-VLM (Compositional Decomposition VLM, Section 3.2.1) and Exec-VLM (Exection VLM, Section 3.2.2). Unlike prior models that encode task information into a single, entangled vector, our architecture explicitly disentangles task comprehension from execution. This decentralized design mirrors the Ventral Stream’s approach to compositional understanding, first compiling knowledge and then acting upon it.
3.2.1 CoDe-VLM: Compiling Tasks into Knowledge Graphs
To build a deep and structured understanding of the task, CoDe-VLM acts as a semantic compiler. It leverages the vast world knowledge embedded within a VLM to on-the-fly compile an unstructured instruction $T$ into a dynamic, task-specific knowledge graph (KG).
CoDe-VLM generates a graph structure encapsulating nodes and relational edges. This task KG, formed from extracted semantic attributes like general description, appearance, and location, constitutes our explicit and compositional representation of the task. This representation not only enables the agent to robustly verify objects encountered during navigation but also provides crucial priors for planning by interfacing with the spatial reasoning components of the Dorsal Stream.
3.2.2 Exec-VLM: Executing Actions via Graph-based Reasoning
The Exec-VLM serves as the agent’s executive core, responsible for determining the optimal action by combining visual observations, spatial awareness from Dorsal Stream, and the structured task semantics provided by CoDe-VLM. Crucially, instead of making decisions in a high-dimensional, entangled feature space, Exec-VLM performs explicit reasoning on the task knowledge graph.
We steer this reasoning process using Chain-of-Thought (CoT). The CoT guides Exec-VLM to break down the complex navigation task into interpretable sub-steps: current state analysis, memory integration, goal analysis, scene assessment, path planning, and action decision. During the “goal analysis” step, for instance, the model directly queries the nodes and edges of the KG to confirm the target’s identity and properties, rather than relying on a fragile memory of the initial instruction.
@@ 3.3 Nav-Ensurance
To enhance the evaluation of safety and efficiency in navigation, we present a new metric Area Overlap Redundancy Index (AORI) (Section 3.3.1). Additionally, we develop Nav-Ensurance, including Multimodal Stuck Detection (Section 3.3.2), context-aware escape strategies (Section 3.3.3), and adaptive precision navigation (Section 3.3.4) to ensure navigation systems reliably and effectively.
@@ 3.3.1 Area Overlap Redundancy Index (AORI)
Where $r_{\text{overlap}}$ represents the ratio of revisited areas, $d_{\text{norm}}$ is the normalized density, and $w_{c}=0.8,w_{d}=0.2$ are weighting coefficients. Further details are provided in Appendix D.
@@ 3.3.2 Multimodal Stuck Detection
To detect if it is stuck, the agent analyzes its trajectory over a sliding window of $T$ steps by computing two key metrics: the progress efficiency $\eta$ and the rotational-to-translational ratio $\rho$.
@@
A stuck state is confirmed if a weighted score $S=w_{\eta}\cdot\mathbb{I}[\eta<\tau_{\eta}]+w_{\rho}\cdot\mathbb{I}[\rho>\tau_{\rho}]$ remains above a threshold $S_{\text{th}}$ for $k$ consecutive windows. These metrics effectively detect situations where the agent makes little forward progress (low $\eta$) or is spinning in place (high $\rho$).
@@ 3.3.3 Context-aware Escape Strategies
When a stuck state is detected, the system selects an appropriate escape strategy based on the perceived information from Dorsal Stream(Section 3.1). For instance, in corner traps (perceived dead ends), a large turn is executed. In narrow passages, a small backward step followed by a randomized direction change is employed. If the environmental context is ambiguous, the agent will analyze recent successful movement directions and attempt to move perpendicularly, significantly improving escape capabilities from complex trap situations.
@@ 3.3.4 Adaptive Precision Navigation
Additionally, when activating the precision navigation mode, the system can utilize visual analysis (using VLM) to create more detailed action options, thereby maximizing final positioning accuracy.
@@ 4 Experiments
We evaluate our proposed DORAEMON within the Habitat simulator [31] on four large-scale datasets: HM3Dv1[28](Object Navigation), HM3Dv2[38](Object Navigation), and MP3D [8](Object Navigation), GOAT[17] (Multi-modal lifelong navigation, using HM3Dv2).
@@
The action space includes stop, move_forward where the distance parameter is sampled from the continuous range $0.5\text{m},1.7\text{m}]$, and rotate. We adopt standard metrics to evaluate navigation performance: Success Rate (SR), the percentage of episodes where the agent successfully stops near a target object; Success weighted by Path Length (SPL), defined as $\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}}{\max(p_{i},l_{i})}$, rewarding both success and efficiency; and our proposed Area Overlap Redundancy Index (AORI) (Equation (3)), which quantifies navigation by penalizing redundant exploration (lower is better). More information is set in the Appendix E.
@@
We compare DORAEMON against several state-of-the-art navigation methods on the HM3Dv2, HM3Dv1, and MP3D. Our main comparison focuses on end-to-end approaches. Beyond these direct end-to-end counterparts, we also consider a broader set of recent methods for non-end-to-end object navigation methods. More baseline details are set in the Appendix I.
@@ 4.1 Methods Comparision
Non-end-to-end methods:
Most methods are non-end-to-end, their reliance on fine-grained discrete actions is a significant departure from natural human behavior, underscoring the superiority of an end-to-end approach. To ensure a fair comparison with these methods that utilize a discrete action set $\mathcal{A}$: move forward 0.25m, turn left/turn right 30, look up/lookdown 30, stop, and a common 500 steps episode limit, we conduct an additional set of experiments. In these, we normalize our agent’s interactions by approximating an equivalent number of standard discrete steps for each of DORAEMON’s actions. During our experiments, one DORAEMON step $t$ was equivalent to about 9-10 non-end-to-end step $t_{n}$.
@@
[…]
Compared to the non-end-to-end approach in the Table 2, DORAEMON achieves state-of-the-art performance on SR, despite normalizing our action to set $\mathcal{A}$. Each action performed by ours corresponds to several actions in this set. In fact, we only run about 60 end-to-end steps, which further demonstrates the excellence of our DORAEMON.(details are provided in the Appendix B)
Ablation Studies.
We perform comprehensive ablation studies to validate our design choices, with results summarized in Table 3. (a) Core Components: Table 3(a) shows that each component is crucial. Removing both the Dorsal and Ventral streams severely degrades performance, confirming their synergistic effect. Disabling the Nav-Ensurance mechanism also notably worsens the AORI, highlighting its effectiveness in error prevention. (b) Choice of VLM: The VLM ablation in Table 3(b) indicates that while Gemini-1.5-Pro is optimal, our framework remains highly effective with smaller models. This demonstrates our architecture’s inherent strength and its plug-and-play nature, suggesting future compatibility with evolving VLMs. (c) Hyperparameter Sensitivity: The analysis in Table 3(c) reveals a trade-off between metrics. For example, setting TopK=12 yields the highest SR but at the cost of SPL. Our default hyperparameters are carefully chosen to achieve a robust and balanced performance across all metrics, rather than over-optimizing for a single one.
Table 3: A comprehensive ablation study of DORAEMON across different datasets, including variations in modules, VLMs, and hyperparameters. All experiments were evaluated over 100 episodes. Our DORAEMON uses the default hyperparameters: TopK=8, memory update interval=3, area grid size=2.
Method / Configuration HM3Dv2 HM3Dv1 MP3D
SR (%) $\uparrow$ SPL (%) $\uparrow$ AORI (%) $\downarrow$ SR (%) $\uparrow$ SPL (%) $\uparrow$ AORI (%) $\downarrow$ SR (%) $\uparrow$ SPL (%) $\uparrow$ AORI (%) $\downarrow$
[…]
w/o Dorsal & Ventral Stream 51.6 18.3 61.5 48.4 18.9 53.7 38.8 13.9 64.3
w/o Dorsal & CoDe-VLM 54.0 19.8 59.1 51.2 19.4 52.5 40.2 14.2 63.8
w/o Dorsal Stream 59.0 22.7 56.3 53.8 20.5 51.1 40.9 14.6 65.1
w/o Nav-Ensurance 60.0 22.5 54.9 53.1 20.7 50.9 42.2 15.3 60.4
[…]
Qwen-7B 49.5 20.6 68.7 - - - - - -
Gemini-1.5-Flash 58.0 20.1 54.8 - - - - - -
Gemini-2-Flash 59.0 21.5 57.9 - - - - - -
[…]
     w/ TopK = 12 65.0 22.57 42.89 - - - - - -
     w/ TopK = 4 59.0 22.78 42.03 - - - - - -
     w/ memory update interval = 1 53.0 19.94 45.54 - - - - - -
     w/ memory update interval = 5 62.0 23.61 44.22 - - - - - -
     w/ area grid size = 1 61.0 21.97 39.66 - - - - - -
     w/ area grid size = 3 61.0 22.95 43.57 - - - - - -
DORAEMON (Ours, default) 61.0 23.7 48.8 55.6 21.4 49.1 41.1 15.8 59.3
4.2 Navigation in Real World
To validate the Sim-to-Real generalization of our model, we deployed our DORAEMON in a novel office environment. Despite the significant domain gap, the agent successfully completed navigation tasks. Figure 5 shows a representative trial. More demos are available on our project homepage.
Figure 5: DORAEMON’s Performance in sim2real
@@ 5 Conclusion
In this paper, we present DORAEMON , a novel cognitive-inspired framework consisting of Ventral and Dorsal Streams that mimics human navigation capabilities. The Dorsal Stream implements the Hierarchical Semantic-Spatial Fusion and Topology Map to handle spatiotemporal discontinuities, while the Ventral Stream combines CoDe-VLM and Exec-VLM to improve decision-making. Our approach also develops Nav-Ensurance to ensure navigation safety and efficiency. Extensive experimental results demonstrate the superior performance of DORAEMON.
@@ Appendix A Action Proposer
DORAEMON employs an Action Proposer[14] to generate a refined set of candidate actions, which the Exec-VLM then evaluates for the final action decision. As shown in Figure 6, first parameterized action candidates $A_{\text{init}}^{t}$ are generated by the parameterized action space (Equation (6)). Second, adaptive filtering (Equation (7)) refines $A_{\text{cand}}^{t}$ using exploration state $\mathcal{V}_{t}$ and historical patterns $\mathcal{H}_{t}$. Safety-critical recovery (Equation(8)) enforces a rotation cooldown $\gamma$ through viability evaluation $\mathcal{F}(\cdot)$. Finally, the projection module visually encodes $A_{\text{final}}^{t}$ into $I^{t}_{\text{anno}}$ with numeric tagging (0 for rotation) to interface with VLM’s semantic space.
Figure 6: Action proposal: (a) Collision-free action generation within $\pm\theta_{\text{max}}$ FOV, (b) Exploration-aware filtering with $\Delta\theta$ angular resolution, (c) Safety-constrained and action projection.
@@ Appendix B Steps Conversion
$T_{\text{episode}}=\sum_{t=1}^{500}t_{n}\leq 500$ (9)
where $t_{n}$ denotes converted steps for action at time step $t$. During our experiments, one DORAEMON step $t$ was equivalent to about 9-10 $t_{n}$
@@
4:  return 11 $\triangleright$ Explicit stop handling
@@
6:  $s_{r}\leftarrow\lceil r/\Delta_{r}\rceil$ $\triangleright$ Radial step calculation
7:  $\theta_{\text{deg}}\leftarrow 180|\theta|/\pi$ $\triangleright$ Radian-degree conversion
8:  $s_{\theta}\leftarrow\lceil\theta_{\text{deg}}/\Delta_{\theta}\rceil$ $\triangleright$ Angular step calculation
9:  $N\leftarrow\max(s_{r}+s_{\theta},1)$ $\triangleright$ Step composition
10:  return NN
@@
We also presented examples of numerical conversions for steps in the experiment.
Table 4: Steps Conversion
(a) Steps Conversion for an end-to-end Step
End-to-end Action Non-end-to-end Steps Non-end-to-end Action
(1.27m, $53^{\circ}$) 9 (1.5m, $75^{\circ}$)/0.25m $\times$ 6 + $25^{\circ}$ $\times$ 3
(1.7m, $60^{\circ}$) 10 (1.75m, $75^{\circ}$)/0.25m $\times$ 7 + $25^{\circ}$ $\times$ 3
(1.1m, $93^{\circ}$) 9 (1.25m, $100^{\circ}$)/0.25m $\times$ 5 + $25^{\circ}$ $\times$ 4
(b) Steps Conversion for a Navigation
End-to-end Steps Non-end-to-end Steps
3 23
16 95
42 300
@@ Appendix C Navigation Case
Figure 8 depicts a full navigation episode in which our memory‑augmented agent searches for a chair in an unfamiliar indoor scene. The seven consecutive frames show the agent’s visual observations and planned motions from entry to target acquisition. Throughout the sequence, the agent (i) reasons about semantic priors—chairs are more likely near tables or in living/dining areas; (ii) fuses transient visual evidence with its episodic memory to avoid revisiting explored regions; and (iii) selects actions that maximise expected information gain while respecting safety constraints. The case illustrates how the proposed memory module complements on‑the‑fly perception to yield efficient, goal‑directed exploration in cluttered, real‑world layouts.
@@
Figure 8: Navigation case Each row shows one decision step. Left: the green circle highlights the action selected for this step. Upper‐right dashed green box displays the most relevant episodic memory retrieved at this step. Lower‐right speech bubble is the agent’s natural‑language rationale that fuses (i) semantic priors, (ii) current visual evidence, and (iii) memory cues.
@@ Appendix D Detailed Description of AORI
D.1 Area Overlap Redundancy Index (AORI)
@@
  • Map resolution: $5,000\times 5,000$ grid (map_size=5000)
@@
  • Optimal Case (AORI=0): $\text{When }r_{\text{overlap}}=0\ \&\ d_{\text{normalized}}=0\Rightarrow 1-(0.8\cdot 1^{2}+0.2\cdot 1)=0$
[…]
@@
[…]
    AORI =1[$0\displaystyle=1-[8(1-1)^{2}+2(1-1)]=0$
@@
[…]